OpenStudy (imstuck):

Hey, can anyone find the pattern hidden in this series of numbers? 3, -6, 12, 4, 20, ? What would be the next number?!!!!

OpenStudy (imstuck):

No one?

OpenStudy (anonymous):

The series is not geometric, and doesn't seem to have enough outputs for wolfram to see a pattern

OpenStudy (imstuck):

I was asked this while I was tutoring on tutor.com and I felt so retarded when I had no answer for the customer. I thought maybe I was going crazy, because I see no discernible pattern at all. Thanks for the confirmation!

OpenStudy (imstuck):

Thanks, also, for taking the time!

OpenStudy (anonymous):

I'm sure you could figure out a pattern, but more terms might be needed

OpenStudy (anonymous):

13

OpenStudy (imstuck):

Gotcha! That's all she posted, so I thought maybe there was an error in her typing, or something!

OpenStudy (imstuck):

Ok, how did you get 13, @OOOPS

OpenStudy (anonymous):

@Opps, I'll leave it to you

OpenStudy (imstuck):

Please enlighten me!

OpenStudy (anonymous):

3-9 =-6, then -6+18= 12 the first couple (-9,+18) 12-8=4 , then 4+16 = 20 (the second couple (-8,16) so the third one should be 20-7=13, then 13+14 =27 to have the third couple is (-7,14) as the series works in previous

OpenStudy (anonymous):

Does it make sense?

OpenStudy (imstuck):

I'm looking at it trying to make sense of it now. Wow! And thanks!

OpenStudy (anonymous):

hihihi if it doesn't, I don't know how to do.

OpenStudy (imstuck):

I can sort of see it, although it's quite mixed up, isn't it! I never in a billion years would have gotten that one! Eeeek!

OpenStudy (imstuck):

Is there a certain method to finding the patterns to numbers? Any great clues or tips?

OpenStudy (anonymous):

just experience, I used to work on repeating series, like 12112111211112..... and I have to figure out the rule. It was nightmare. hihihi...